OpenStudy (anonymous):

HELP PEOPLE HARD MATH QUESTION :0

OpenStudy (anonymous):

Simplify the expression \[\sqrt{27x ^{5}} + x \sqrt{3} + 9\sqrt{3x}\] If the simplified expression is written in standard form, what is the leading coefficient?

OpenStudy (across):

Hint:\[3^3=27\]

OpenStudy (anonymous):

so the answer is 3 lol?

OpenStudy (anonymous):

but it's a sqrt not a cubed root?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

across stated that 3cubed is 27 i'm waiting for her to reply to help us understand why o.o

OpenStudy (anonymous):

oh yea ^^

OpenStudy (anonymous):

the only coefficient i see is 2

OpenStudy (anonymous):

across is taking forever to answer lol

OpenStudy (anonymous):

wait did we get the answer?

OpenStudy (anonymous):

you just deleted that whole thing?!

OpenStudy (unklerhaukus):

i made an error

OpenStudy (anonymous):

what ^^

OpenStudy (unklerhaukus):

\[\quad\sqrt{27x ^{5}} + x \sqrt{3} + 9\sqrt{3x}\]\[=\sqrt{3^3}x ^{5/2} + \sqrt{3}x^1 + 9\sqrt{3}x^{1/2}\]\[=3\sqrt{3}x^2x ^{1/2} + \sqrt{3}x^{1/2}x^{1/2} + 9\sqrt{3}x^{1/2}\]\[=(x^2 + x^{1/2} + 9)\sqrt 3x^{1/2}\]\[=(x^2 + x^{1/2} + 9)\sqrt {3x}\]

OpenStudy (anonymous):

so thats the answer LOL

OpenStudy (unklerhaukus):

hmm

OpenStudy (anonymous):

the last part is right?

OpenStudy (unklerhaukus):

maybe my second line in more useful

OpenStudy (anonymous):

what do u mean?